OpenStudy (anonymous):

Find the value of the variable(s). If your answer is not an integer, leave it in simplest radical form.

OpenStudy (anonymous):



OpenStudy (anonymous):

A. \[\frac{ 1 }{ 2 }\] B. \[16\sqrt{3}\] C. 2 D. \[8\sqrt{3} \]

OpenStudy (ranga):

You can use the definition of sine to find x. You can also use Pythagoras theorem: a^2 = b^2 + c^2 to find a.

OpenStudy (ranga):

to find x.

OpenStudy (anonymous):

ratios of a 30 - 60 - 90 right triangle are \[1:\sqrt3:2\]

OpenStudy (anonymous):

I did pythagorean theorem and got 320=x^2 I dont know which answer choice = \[\sqrt{320}\] ....

OpenStudy (ranga):

you added the two squares. you should subtract. 16^2 = 8^2 + x^2 x^2 = 16^2 - 8^2

OpenStudy (anonymous):

So...its \[\sqrt{192}\] and that is equivalent to 8 \[\sqrt{3} \]. is the answer D?